On April 1,2020 John’s construction company purchased equipment costing 68,000 the equipment is estimated to have a nine year useful life with no salvage value calculate the amount of the adjustment at the December 31, 2020 year end


1
Expert's answer
2021-02-28T11:37:56-0500

The Estimated Useful Life (denoted by EUL) of equipment is equal to 9 years or 108 months.

Let's compute the depreciation expense per month (denoted by DM) using the Straight-line method. This method is based on the following formula:

"DM=(IEP-SV)\/EUL"

where IEP is initial equipment price and SV is a salvage value.

Hence,

"MD=(68,000-0)\/108=629.63"

Amount of the adjustment at the December 31, 2020 year end (accumulated deprecation) can be calculated as following

"T\times MD=9*629.63=5,666"

The Carrying Value (adjusted value) of equipment at the December 31, 2020 is 62,333 (calculated as 68,000 - 62,333).